OpenStudy (anonymous):

hw to find (d^2y/dx^2) if (2x^3-3y^2=8)

OpenStudy (asnaseer):

Use implicit differentiation. This video might help if you are unfamiliar with the technique: http://www.youtube.com/watch?v=5yTVUZCaU6k

OpenStudy (anonymous):

my answer is 2x

OpenStudy (asnaseer):

how did you get that - please show your steps

OpenStudy (anonymous):

\[2x^3-3y^2=\]

OpenStudy (anonymous):

\[6x^2-6y \frac{ dy }{ dx}=0\] \[12x-6\frac{ d^2y }{dx^2 }=0\] \[6\frac{ d^2y }{dx^2 }=12x\] \[\frac{ d^2y }{ dx^2}=2x\]

OpenStudy (asnaseer):

the second term in the second step above is incorrect. you need to use the product rule here.

OpenStudy (anonymous):

how to use the product rule.. it mean udv+vdu>>> is this the formula.

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

you might find it easier to write y' and y'' instead of dy/dx etc

OpenStudy (anonymous):

which part must be as u and v..

OpenStudy (asnaseer):

it doesn't matter - the rule is symmetric :)

OpenStudy (asnaseer):

d(uv)/dx = d(vu)/dx

OpenStudy (amistre64):

might be easier to see the chain rule and product rule if y is written in function notation y(x) just a thought. It tends to make the equations bulkier, but easier to apply rules to in the learning process

OpenStudy (asnaseer):

-6y*y' is the term that you need to apply the product rule to. The -6 is just a constant multiplier so you can ignore that and just multiply the result by -6 at the end.

OpenStudy (amistre64):

\[2x^3-3y^2=8\] \[2x^3-3(y(x))^2=8\] \[6x^2-6y(x)y'(x)=0\] etc

OpenStudy (anonymous):

ok if i applied for u=-6 v=y so the answer should be -6(dy/dx)+0----- using product rule.

OpenStudy (asnaseer):

no - the -6 is just a constant. as I said above, you need to apply it to y*y'. so, for example, let u=y and let v=y'

OpenStudy (asnaseer):

whatever answer you get, multiply it all by -6 at the end

OpenStudy (anonymous):

ok..i thought that u=-6.. sorry for the mistake.

OpenStudy (asnaseer):

np :)

OpenStudy (asnaseer):

alternatively, you can take u=-6y and v=y'

OpenStudy (anonymous):

so now i get \[12x-6(y \frac{ d^2y }{ dx^2 }+\frac{ d^2y }{ dx^2 }))\] so can i solve this problem

OpenStudy (asnaseer):

still not correct. try and do just the derivative of this first:\[-6y\frac{dy}{dx}\]

OpenStudy (anonymous):

wait first (dy/dx*dy/dx)=d^2y/dx^2

OpenStudy (asnaseer):

no - y' * y' = (y')^2

OpenStudy (asnaseer):

(y')^2 is not the same as y''

OpenStudy (anonymous):

ok.. for -6y(dy/dx)= 6y(y'')-6(y')^2

OpenStudy (asnaseer):

yes - that looks correct now.

OpenStudy (asnaseer):

wait - the initial term should also be negative

OpenStudy (anonymous):

sorry i fogot to put. so how i going to solve it.

OpenStudy (asnaseer):

write out the final expression

OpenStudy (anonymous):

12x-6y(y'')-6(y')^2

OpenStudy (asnaseer):

you are missing the equals sign and the value on the right-hand-side

OpenStudy (anonymous):

12x-6y(y'')-6(y')^2=0

OpenStudy (asnaseer):

good - so first thing to spot is that you can divide through by 6

OpenStudy (anonymous):

2x-y(y'')-(y')^2=0 so how i going to get y''

OpenStudy (asnaseer):

now rearrange this to solve for y''. you can substitute in the value for y' from your earlier step:\[6x^2-6y \frac{ dy }{ dx}=0\]

OpenStudy (anonymous):

y'=x^2/y

OpenStudy (asnaseer):

correct. use this in your equation involving y'' and then just rearrange to get y''

OpenStudy (anonymous):

final answer = y''=x^4/y^3-(2x/y)

OpenStudy (asnaseer):

doesn't look quite right

OpenStudy (asnaseer):

still not correct

OpenStudy (anonymous):

sorry y''=2x/y-(x^4/y^3)

OpenStudy (asnaseer):

nope

OpenStudy (asnaseer):

oops - sorry - misread - it IS correct

OpenStudy (amistre64):

maybe easier to work thru this way, just an idea \[D_x[2x^3-3y^2=8]\] \[6x^2-6y~y'=0\] \[-6y~y'=-6x^2\] \[y'=\frac{x^2}{y}\] \[y'=x^2y^{-1}\] \[D_x[y'=x^2y^{-1}]=\frac{d^2y}{dx^2}\]

OpenStudy (asnaseer):

nice @amistre64 :)

OpenStudy (anonymous):

thank you very much.. all

OpenStudy (asnaseer):

yw :)